OpenStudy (anonymous):

For a summer fundraiser, Rob is doing manual labor to help several families do home improvement projects. He is tracking the amount of time he spends each week with a spreadsheet. If his spreadsheet shows that during the past two weeks, he spent a total of six thousand minutes working, how many days and hours has he worked? 1 day and 7 hours 1 day and sixteen hours 4 days and 2 hours 4 days and 4 hours

OpenStudy (anonymous):

I'm doing this on a library computer and I only have about twenty minutes left.... I wouldn't ask, except I am completely lost....

OpenStudy (loser66):

\[\dfrac{1day}{24hours}\\\dfrac{1hour}{60minutes}\] so, just apply \[\dfrac{1day}{24hours}*\dfrac{1hour}{60minutes}*6000(minutes) \] to get the answer

OpenStudy (anonymous):

Thx, I wasn't sure what to do

OpenStudy (anonymous):

24 hrs. x 60 mins. x 6000 mins.?

OpenStudy (loser66):



OpenStudy (anonymous):

okay, I get it now, thx

OpenStudy (anonymous):

can I ask one more Q?

OpenStudy (anonymous):

A line segment with endpoints Q negative six comma nine and R twelve comma negative three is dilated with the origin as the center of dilation. The resulting endpoints are Q prime negative two comma three and R prime four comma negative one. What is the scale factor? one-ninth one-third 3 9

OpenStudy (anonymous):

The first one you helped me with I got right!! THX!! @Loser66